OpenStudy (anonymous):

You purchase 10,000 shares of a stock for $0.85 a share. One year later it is worth $0.95 a share. Find the percent of increase

OpenStudy (anonymous):

11.76 %

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

It increased from 85/100 to 95/100. So increase is 10/100. Increase % = (Increase/initial amount)*100 = (10/100) / (85/100) * 100 = 200/17 Divide it and it would be approximately 11.76%.

OpenStudy (anonymous):

ty